A. You are provided with the following information form the accounts of BBS Ltd for the year ending 30 June 2019 Cash Sales 950 000 Cost of Goods Sold 35 000 Amount received in advance for services to be performed in August 2019 9 500 Rent expenses for year ended 30 June 2019 9 000 Rent Prepaid for two months to 31 August 2019 1 200 Doubtful debts expenses 1 200 Amount provided in 2019 for employees’ long-service leave entitlements 5 000 Goodwill impairment expenses 7 000 Required: Calculate the taxable profit and accounting profit for the year ending 30 June 2019. B. GYV Ltd has the following deferred tax balances as at 30 June 2019. Deferred tax asset $9 00 000 Deferred tax liability $7 00 000 The above balances were calculated when the tax rate, was 20 percent. On 1 December 2019, the government raises the corporate tax rate to 25 percent. Required: Provide the journal entries to adjust the carry-forward balances of the deferred tax asset and deferred tax liability
Answer #1

ANSWER:

(1).

Calculate taxable profit and accounting profit for the year ending 30 June 2019

Calculate the accounting profit:
Particulars Amount ($)
Cash sales $950,000
Less: Cost of Goods sold $35,000
Gross profit $915,000
Less: Expenses
Rent expense $9,000
Doubtful debts $1,200
Amount paid for Employees LSL (50% upto June, 2019) $2,500
Goodwill impairment expense $7,000
Total accounts profit $895,300

..

Taxable profit

Particulars

Amount ($)

Total accounts profit

$895,300

Less: Amount paid for Employees LSL (50%)

2500

Add : amount received in advance

9500

Less: Prepaid rent

1200

Add: doubtful debt exp

1200

Total Taxable profit

$902300

(2).

Changes In Future Tax Rates

20% rate to 25%

Here the Tax rate increase:

DTA will increase

Current deferred tax assets based on 20% =900000

If rate change to 25% = DTA = 900000 / 20% * 25% = 1125000

The DTA increased by $225000 ( 1125000 - 900000 )

Journal entry

Accounts

Debit

Credit

Deferred Tax Asset

$225000

Income Tax Expense

$225000

(To record adjustment in tax rate)

.

DTL will decrease

Current deferred tax Liability based on 20% =700000

If rate change to 25% = DTA = 700000 / 20% * 25% = 875000

The DTL increased by $175000 ( 875000 - 700000 )

Journal entry

Accounts

Debit

Credit

Income Tax Expense

$175000

Deferred Tax Liability

$175000

(To record adjustment in tax rate)

=================

DEAR STUDENT,

IF YOU HAVE ANY QUERY PLEASE ASK ME IN THE COMMENT BOX,I AM HERE TO HELP YOU.PLEASE GIVE ME POSITIVE RATING..

****************THANK YOU******************